The lengths of two sides or a triangle are 7 and 10. The length of the third side may be:
a)17 b)20 c)3 d)8

Answer by venugopalramana(3286)

THE RULE IS THAT IN A TRIANGLE , SUM OF ANY 2 SIDES SHOULD BE GREATER THAN THE THIRD SIDE.SO THE ANSWER IS D,THAT IS 8,SINCE WITH,7,10 AND 8 WE HAVE
7+10>8,7+8>10 AND 8+10>7
A)17 IS NOT CORRECT BECAUSE 7+10=17...SUM OF 2 SIDES IS NOT GREATER THAN THE THIRD SIDE
B)20 IS NOT CORRECT BECAUSE 7+10=17<20....SUM OF 2 SIDES IS NOT GREATER THAN THE THIRD SIDE
C)3 IS NOT CORRECT BECAUSE 7+3=10... SUM OF 2 SIDES IS NOT GREATER THAN THE THIRD SIDE